The point (6,-3) is on the graph of y=f(x). Find the corresponding point on the graph of g(x)=f(3x)
Big troubles on this one for some reason...any help is greatly appreciated.

Answer by richard1234(7193)

All we know is f(6) = -3 and g(x) = f(3x). The only point on g(x) that we can find occurs when x = 2. This is because g(2) = f(3*2) = f(6) (in other words, I basically set 3x = 6). Since we know f(6) = -3, g(2) must also be -3, so (2,-3) is the corresponding point.